OpenStudy (anonymous):

what is ds/dt if t/(9t+8)

OpenStudy (anonymous):

The answer is potatoe.

zepdrix (zepdrix):

Recall the quotient rule porsche? :)

OpenStudy (anonymous):

for real? comeon bro

OpenStudy (anonymous):

its a derivitave

zepdrix (zepdrix):

yes its a method for taking the derivative :O

OpenStudy (anonymous):

how do i do it?

zepdrix (zepdrix):

\[\frac{ d }{ dt }(\frac{ u }{ v })=\frac{ u'v-uv' }{ v^2 }\] So in your example, t=u, 9t+8=v :D

OpenStudy (anonymous):

im legit so lost

OpenStudy (anonymous):

ohhh shi

OpenStudy (anonymous):

i got it, i think

OpenStudy (anonymous):

nvm im still fluttering lost

zepdrix (zepdrix):